1.27um emission from Venus night side on Dec 27, 2010 1 The following spectra are of the night sky. The flux is negative due to a slight oversubtraction of the telluric spectrum because the sky flux was so low. THe wavelength range is from 1.2-1.3 um and no discernible emission at 1.27um is detected.